SOME DEFINITION OF THE ARTIN EXPONENT OF FINITE GROUPS

K. K. NWABUEZE

Abstract. The Artin exponent induced from cyclic subgroups of finite groups was studied extensively by T.Y. Lam. A Burnside ring theoretic version of Lam’s results for p-groups was given by the author in an earlier paper. Here we look at the Artin exponent induced from the elementary abelian subgroups of finite p-groups using some results of A. Dress.